OpenStudy (anonymous):

Aaaaand here goes the rest of the front sheet. 2. Find the slopes of the lines AB and CD. Are they parallel? a. A (0,-6) B (4,-4) C (3,2) D (2,3) b. A (2,6) B (4,7) C (4,-5) D (8,-2)

OpenStudy (aroub):

a) AB = 1/2 CD = -1 b)AB=1/2 CD=-3/4 No,they are not parallel.